OpenStudy (anonymous):

Solve for x, X^2+6x-40=0

OpenStudy (anonymous):

You can either factor or use the quadratic formula.

OpenStudy (anonymous):

I am not quite sure how to do that.

OpenStudy (anonymous):

This one is factorable, so lets do that.

OpenStudy (anonymous):

Because it is an x^2 with no number multiplied by it, we know it will be in the form (x+?)(x+?)

OpenStudy (anonymous):

Have you done FOIL yet?

OpenStudy (anonymous):

No I have no idea what that is

OpenStudy (anonymous):

If we have something like this (x-4)(x+3) you can expand it by using a method called FOIL.

OpenStudy (anonymous):

First Outside Inside Last

OpenStudy (anonymous):

you multiply the terms like that and add them together

OpenStudy (anonymous):

in this example, we would end up with x^2 +3x -4x -12

OpenStudy (anonymous):

which simplifies to x^2 -x -12

OpenStudy (anonymous):

Now, notice that that looks a whole lot like the problem you started with! We are going to do FOIL backwards.

OpenStudy (anonymous):

X^2+6x-40 we are looking for two numbers that when multiplied give us -40, but when added give us 6.

OpenStudy (anonymous):

they can be positive or negative, and they'll be integers in a basic problem such as this. Any numbers come to mind?

OpenStudy (anonymous):

How about 4 and -10?

OpenStudy (anonymous):

So, we end up with this: (x+4)(x-10) When FOILed, we get this: X^2+6x-40 Remember, this is set equal to zero so we have ended up with (x+4)(x-10)=0

OpenStudy (anonymous):

Now we can solve it. This equals zero when x is either -4 or 10.

OpenStudy (anonymous):

This is how I got that: you can set both parts equal to zero: x-4=0 x+10=0 solving, we get x to equal -4 and 10.

OpenStudy (anonymous):

oops, i meant this: This is how I got that: you can set both parts equal to zero: x+4=0 x-10=0 solving, we get x to equal -4 and 10. Quadratic equations like these have two answers the majority of the time, so the answer would be x=-4, 10

OpenStudy (anonymous):

err, scratch that

OpenStudy (anonymous):

Back when I first said "How about 4 and -10?" I got the signs backwards. So x actually comes out to equal 4 and -10 in the end... sorry

OpenStudy (anonymous):

Did you get all that? Kinda confusing the first time around...

OpenStudy (anonymous):

It's fine. Thank you for your help! I really appreciate it. I kind of understand. Im pretty bad at math.